OpenStudy (anonymous):

Evaluate the following integral: (4-2x)dx [2,5] I can get it part way then I freeze.

OpenStudy (nubeer):

\[\int\limits_{2}^{5} 4 -\int\limits_{2}^{5} 2x\] can you do now?

OpenStudy (anonymous):

\[\int\limits_{2}^{5} 4(dx) = 4x - 2\int\limits_{2}^{5} x(dx) = x^{2}?????\] that is where I get up to and now I am stuck

OpenStudy (nubeer):

hmm let me remind u the rule ∫x^n = [x^(n+1)]/ (n+1)

OpenStudy (nubeer):

there is 1 added in the power.. and its divided by the new power we got.

OpenStudy (anonymous):

so x^2/2 right?

OpenStudy (nubeer):

yes. :)

OpenStudy (anonymous):

So do I then do (5-2)^2/2?

OpenStudy (nubeer):

hmm? how you got 5 here? its like 4x - x^2

OpenStudy (nubeer):

and aplly the limit on this.

OpenStudy (anonymous):

so 4(5-2) - (5-2)^2?

OpenStudy (anonymous):

nope, wrong answer.

OpenStudy (anonymous):

the limit is [2,5] right?

OpenStudy (nubeer):

hmm no.. that is not how we apply limit.. [4(5)-(5)^2] - [ 4(2)-(2)^2]

OpenStudy (anonymous):

ahhhhhh ..... grrrrrr. so this equals -9.

OpenStudy (nubeer):

yes . i think so.. is it wrong?

OpenStudy (anonymous):

nope that is what the back of the book says. What ever happened to that 2 that was before the integral?

OpenStudy (nubeer):

hmm no look